OpenStudy (anthonyn2121):

Find dy/ds if y= s sqrt(1-s^2)+arccos(s)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=s\sqrt{1-s^2}+\cos^{-1}s }\) What problem do you have with differentiating the above?

OpenStudy (solomonzelman):

Can you use the *Product Rule* for the following? \(\large\color{black}{ \displaystyle s\sqrt{1-s^2} }\)

OpenStudy (anthonyn2121):

I used the product rule and derived the arccos and got